# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ III.1 Hawking radiation
$T_{H}=\frac{\hbar\kappa_{+}}{2\pi}=\frac{\hbar}{8\pi M}$ (III.1)
@@
$L=\frac{\pi^{2}}{60\hbar^{3}}16\pi M^{2}\frac{\hbar^{4}}{8^{4}\pi^{4}M^{4}}=\frac{\hbar}{15360\pi M^{2}}$ (III.3)
@@
$L=\frac{\hbar}{M^{2}}P$ (III.4)
where $P$ is some dimensionless constant, which needs to be computed numerically. According to Page [27], the constant is equal to $P=3.366\times 10^{-5}$ for photon emissions, and $P=0.3845\times 10^{-5}$ for graviton emissions, and according to Elster [28] $P=7.44\times 10^{-5}$ for scalar field.
In the case of RN black hole, the luminosity depends also on the charge, and it is equal to:
$L=\frac{\hbar}{M^{2}}H_{M}\left(\frac{Q}{M}\right)$ (III.5)
for some dimensionless function $H_{M}$ (that may be determined numerically).
@@
$M_{,v}^{(+)}=M_{,t}^{(\infty)}=-L=-\frac{\hbar}{M^{2}}H_{M}\left(\frac{Q}{M}\right)$ (III.7)
@@
$M(v)=\left(M_{0}^{3}-3v\hbar H_{M}\left(\frac{Q}{M}\right)\right)^{\frac{1}{3}}$ (III.8)
@@
$M(v)=\left(M_{0}^{3}-M_{0}^{2}\varepsilon_{0}v\right)^{\frac{1}{3}}$ (III.9)
@@
$M_{,v}=-\frac{1}{3}M_{0}^{2}\varepsilon_{0}\left(M_{0}^{3}-M_{0}^{2}\varepsilon_{0}v\right)^{-\frac{2}{3}}$ (III.10)
@@
$T_{ev}=\frac{M_{0}^{3}}{3\hbar H_{M}\left(\frac{Q}{M}\right)}$ (III.12)
@@
$\frac{T_{ev}}{M_{0}}=\frac{M_{0}^{2}}{3\hbar H_{M}\left(\frac{Q}{M}\right)}=\frac{1}{\varepsilon_{0}}$ (III.13)
III.2 Choice of initial mass parameter
Since we wish to cover significant evaporation in our numerical simulation (we typically manage to evaporate about 53% of the initial mass), our numerical grid must cover a grid length of order $\frac{T_{ev}}{M_{0}}\left(\frac{du}{M_{0}}\right)^{-1}$points. ($du$ is the size of a single grid step, and we usually have $\frac{du}{M_{0}}=\frac{1}{100}$.) For astrophysical black holes of 10 solar masses, $\frac{T_{ev}}{M_{0}}\approx 10^{82}$, which is impossible to implement numerically. We can implement numerically $\frac{T_{ev}}{M_{0}}=10^{3}$. This value definitely does not correspond to a realistic astrophysical black hole, yet this value is still $\gg 1$. We’ll therefore choose our $M_{0}$ such that $\frac{T_{ev}}{M_{0}}=10^{3}$. The corresponding value of $\varepsilon_{0}$ is $\frac{1}{1000}$.
III.3 Choice of units
@@
Recall the parameter $L_{0}$ from Eq. (II.1), which can be chosen freely. We choose $L_{0}=M_{0}$. In our choice of units, we therefore get $L_{0}=1$.
III.4 Basic scaling of semiclassical $T_{\alpha}^{\;\beta}$
@@
$T_{\alpha}^{\;\beta(SC)}=\frac{\hbar}{M^{2}}\hat{T}_{\alpha}^{\;\beta}$ (III.14)
@@
$T_{\alpha}^{\;\beta(SC)}=\frac{\hbar}{M^{4}}H_{T_{\alpha}^{\;\beta}}\left(\frac{r}{M},\frac{Q}{M}\right)$ (III.15)
@@
$T_{vv}\equiv T_{vv}^{(SC)}=\frac{\hbar}{M^{4}}H_{T_{vv}}\left(\frac{r}{M},\frac{Q}{M}\right)$ (III.16)
@@
In Eq. (II.43) we defined the quantity $\widetilde{T}_{vv}\equiv 8\pi r^{2}T_{vv}$. Correspondingly we define $\widetilde{T}_{vv}^{(+)}$ , $\widetilde{T}_{vv}^{(-)}$ , which are $\widetilde{T}_{vv}$ at the two horizons, EH and IH respectively. They scale according to:
$\widetilde{T}_{vv}^{(+)}=8\pi r_{+}^{2}\frac{\hbar}{M^{4}}H_{T_{vv}}\left(\frac{r_{+}}{M},\frac{Q}{M}\right)\equiv\frac{\hbar}{M^{2}}H_{T_{vv}}^{(+)}\left(\frac{Q}{M}\right)$ (III.17)
$\widetilde{T}_{vv}^{(-)}=8\pi r_{-}^{2}\frac{\hbar}{M^{4}}H_{T_{vv}}\left(\frac{r_{-}}{M},\frac{Q}{M}\right)\equiv\frac{\hbar}{M^{2}}H_{T_{vv}}^{(-)}\left(\frac{Q}{M}\right)$ (III.18)
III.5 The charge function $Q(u,v)$
We would like the charge of the black hole to evaporate as well as the mass, so that the mass could evaporate completely. (If the charge does not evaporate, the mass will stop evaporating when the black hole approaches extremality.) If there is a charged quantum field, then $Q$ will indeed evaporate, and we will have electric currents in spacetime. We will therefore assume that there is such a charged quantum field. Hence $Q$ is no longer a constant parameter, but a function $Q(u,v)$ of $u$ and $v$. At the horizon, we have
$Q=Q^{v}(v)$ (III.19)
@@ III.5 The charge function $Q(u,v)$
$Q(u,v)=Q^{v}(v)+\delta Q(u,v)$ (III.20)
Note that from Eq. (II.28) if $J_{u}=0$ throughout, then $Q_{,u}=0$, and hence $\delta Q(u,v)=0$ because $\delta Q$ vanishes at the EH.
@@
$\frac{Q^{v}(v)}{M(v)}=const$ (III.21)
To what extent is this choice realistic? Let us consider the (arguably hypothetical  [^3] This case is not regarded as realistic because the plasma that typically surrounds black holes would discharge them very quickly.) case of astrophysically large black hole which is significantly charged initially. Such a black hole would lose charge due to the Schwinger pair production effect, where the strong electric field near the EH causes electron-positron pair to be created from the vacuum: one of the charged particles would fall into the black hole, reducing its charge, and the other would escape to infinity. This effect, if possible, would quickly discharge the black hole, especially since the charge-to-mass ratio for electrons is $\approx 2\cdot 10^{21}$. But Schwinger pair production is effective only if the electric field at the event horizon exceeds a critical value $E_{cr}$. The electric field at the EH near extremality is of order $\frac{Q}{r_{+}^{2}}\approx\frac{Q}{M^{2}}=\frac{Q}{M}\frac{1}{M}$. Therefore, if $\frac{Q}{M}\frac{1}{M}>E_{cr}$, Schwinger pair production will quickly discharge the black hole until $\frac{Q}{M}\frac{1}{M}=E_{cr}$. This means there is a critical mass $M=\frac{1}{E_{cr}}$, where a black hole more massive than this critical mass will discharge very slowly (as Schwinger pair production is greatly reduced beyond this critical mass). In this range the black hole would lose its mass until it reaches extremality. On the other hand, a black hole below this critical mass will discharge while approximately keeping $\frac{Q}{M}=ME_{cr}$. This means that in this range, $\frac{Q}{M}$ decreases linearly with $M$. This varying $\frac{Q}{M}$ would result in a more complicated dynamics inside the evaporating black hole, which we will here avoid for the sake of simplicity.
@@
$\delta Q=\log\left(1+e^{\frac{u+v}{M(v)}}\right)M(v)^{-1}\varepsilon_{0}M_{0}^{2}\alpha$ (III.22)
This structure of $\delta Q$ contributes currents flowing in the $r_{*}=\frac{u+v}{2}$ direction (at leading order in $\varepsilon_{0}$). $\alpha$ is a free dimensionless parameter controlling the strength of the semiclassical currents, which we typically choose to be of order unity. For the dependence of $\delta Q$ on $u+v$, we chose it to be via the function $f(x)=\log\left(1+e^{x}\right)$ as it interpolates smoothly between $f(x)=0$ at $x\ll 0$ (as desired at the EH) to $f(x)=x$ at $x\gg 0$ (corresponding to region 3). The argument of this function $f(x)$ is $x=\frac{u+v}{M(v)}$, which is a dimensionless variable proportional to $r_{*}=\frac{u+v}{2}$. Since $\delta Q$ is initiated from semiclassical currents, it should be proportional to $\hbar$ ($\propto\varepsilon_{0}M_{0}^{2}$). The later has mass squared dimensions, and the dimensions of $Q$ is $M^{1}$, therefore we also need to multiply the expression by $M(v)^{-1}$ to get the correct mass dimensions.
The main motivation for introducing $\delta Q$ is to demonstrate numerically that semiclassical drift in the charge $Q$ that may take place and even become significant in region 3 does not affect the backreaction of the metric in region 3 – once the four input functions $M(v)$, $Q^{v}(v)$, $\widetilde{T}_{uu}^{(-)}(v)$, $\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ have been specified. (Note that $\delta Q$ may have an effect on the value of $\widetilde{T}_{uu}^{(-)}(v)$, $\text{$\widetilde{T}_{vv}^{(-)}$}(v)$, similar to that of the sources $Z_{R}$$Z_{S}$.) This fact is already embodied in the analytical approximation – where the charge function used is $Q^{v}(v)$ only. Note that in the absence of $\delta Q$, from Eq. (II.28) $J_{u}=0$. Non-vanishing $\delta Q$ allows us to explore scenarios with $J_{u}$ as well.
III.6 Choice of source functions $Z_{R}$, $Z_{S}$
As already mentioned in subsection II.8, for the task of validating the analytical approximation, we can freely choose the semiclassical functions which are $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$. This amounts to freely choosing the two source functions $Z_{R}$, $Z_{S}$ defined in Eqs. (II.119II.120).
@@ III.6 Choice of source functions $Z_{R}$, $Z_{S}$
$Z_{R}=\varepsilon_{0}M_{0}^{2}M^{-3}e^{S}L_{0}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)$ (III.23)
$Z_{S}=\varepsilon_{0}M_{0}^{2}M^{-5}e^{S}L_{0}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)$ (III.24)
@@
The reasoning behind this choice is as follows. In the semiclassical theory these source functions vanish both at the EH and the IH. We therefore choose these functions to be proportional to $e^{S}$, which is close to zero near both the EH and IH [^4] For RN background in Unruh state, it was found [32] that near the IH $Z_{R}$ is proportional to $e^{S}\sim e^{-2\kappa_{-}r_{*}}$, whereas $Z_{S}$ decays as $r_{*}^{-5}$. In our numerical simulation, we decided to simplify the computation scheme by assuming both to be proportional to $e^{S}$ near the IH. We do not expect this to change the resultant backreaction metric because $r_{*}^{-5}$ vanishes sufficiently fast.. We then multiply $e^{S}$ by the length scale $L_{0}$ similar to the way it appears in the evolution equations, see Eqs. (II.38,II.39). The source term is semiclassical, therefore it should be proportional to $\hbar$ ($\propto\varepsilon_{0}M_{0}^{2}$) which has mass squared dimensions, so we want to create a dimensionless quantity from that, leading us to the dimensionless factor $\hbar M^{-2}\propto\varepsilon_{0}M_{0}^{2}M^{-2}$. Since $F_{1}(R)]=\left[\frac{1}{\sqrt{R}}\right]=[l]^{-1}$, the $Z_{R}$ source term is overall multiplied by $M^{-3}$, and since $F_{2}(R)]=\left[\frac{1}{R^{\frac{3}{2}}}\right]=[l]^{-3}$, the $Z_{S}$ source term is overall multiplied by $M^{-5}$. For convenience, we chose the spacetime dependence of these sources to be dependent on $\frac{r}{M}$ only, and we further restricted our choice to linear functions of this variable, determined by the four arbitrary parameters $Z_{R0}$$Z_{Rr}$$Z_{S0}$$Z_{Sr}$. When running our numerical code, we’ll choose multiple specific (arbitrary) combinations of these parameters and the charge parameter $\alpha$.
@@
The presence of these source terms, and / or the changing charge, cause non-conservation of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$, as already mentioned in subsection II.8. We shall now calculate the amount of this non-conservation, using the evolution equations and the source terms in Eqs. (III.23,III.24). We start by taking the derivative of the constraint Eq. (II.44):
$\left(\widetilde{T}_{uu}\right)_{,v}$ $=R_{,uv}S_{,u}+R_{,u}S_{,uv}-\left(R_{,uv}\right)_{,u}$ (III.25)
$=L_{0}e^{S}\left(F_{1}(R)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)S_{,u}$ (III.26)
$+\left(e^{S}L_{0}F_{2}(R)+\varepsilon_{0}M^{-5}e^{S}L_{0}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,u}$ (III.27)
$-\left[L_{0}e^{S}\left(F_{1}(R)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)\right]_{,u}$ (III.28)
$=L_{0}e^{S}\left(F_{1}(R)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)S_{,u}$ (III.29)
$+\left(e^{S}L_{0}F_{2}(R)+\varepsilon_{0}M^{-5}e^{S}L_{0}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,u}$ (III.30)
$-L_{0}e^{S}\left(F_{1}(R)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)S_{,u}$ (III.31)
$-e^{S}L_{0}\frac{dF_{1}(R)}{dR}R_{,u}-e^{S}L_{0}\frac{dF_{1}(Q,R)}{dQ}Q_{,u}-\varepsilon_{0}M^{-3}e^{S}L_{0}\frac{r_{,u}}{M}Z_{Rr}$ (III.32)
$=e^{S}L_{0}\left[\left(\varepsilon_{0}M^{-5}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,u}-\frac{2Q}{R^{\frac{3}{2}}}Q_{,u}-\varepsilon_{0}M^{-3}\frac{r_{,u}}{M}Z_{Rr}\right]$ (III.33)

### v2

@@ 3.1 Hawking radiation
$T_{H}=\frac{\hbar\kappa_{+}}{2\pi}=\frac{\hbar}{8\pi M}$ (3.1)
@@
$L=\frac{\pi^{2}}{60\hbar^{3}}16\pi M^{2}\frac{\hbar^{4}}{8^{4}\pi^{4}M^{4}}=\frac{\hbar}{15360\pi M^{2}}$ (3.3)
@@
$L=\frac{\hbar}{M^{2}}P$ (3.4)
where $P$ is some dimensionless constant, which needs to be computed numerically. According to Page [27], the constant is equal to $P=3.366\times 10^{-5}$ for photon emissions, and $P=0.3845\times 10^{-5}$ for graviton emissions, and according to Elster [28] $P=7.44\times 10^{-5}$ for scalar field. There are two degrees of freedom for the photon and graviton while there is a single degree of freedom for the scalar.
In the case of RN black hole, the luminosity depends also on the charge, and it is of the form:
$L=\frac{\hbar}{M^{2}}H_{M}\left(\frac{Q}{M}\right)$ (3.5)
for some dimensionless function $H_{M}\left(\frac{Q}{M}\right)$ of the variable $\frac{Q}{M}$ (that may be determined numerically).
@@
$M_{,v}^{(+)}=M_{,t}^{(\infty)}=-L=-\frac{\hbar}{M^{2}}H_{M}\left(\frac{Q}{M}\right)$ (3.7)
@@
$M(v)=\left(M_{0}^{3}-3v\hbar H_{M}\left(\frac{Q}{M}\right)\right)^{\frac{1}{3}}$ (3.8)
@@
$M(v)=\left(M_{0}^{3}-M_{0}^{2}\varepsilon_{0}v\right)^{\frac{1}{3}}$ (3.9)
@@
$M_{,v}=-\frac{1}{3}M_{0}^{2}\varepsilon_{0}\left(M_{0}^{3}-M_{0}^{2}\varepsilon_{0}v\right)^{-\frac{2}{3}}$ (3.10)
@@
$T_{ev}=\frac{M_{0}^{3}}{3\hbar H_{M}\left(\frac{Q}{M}\right)}$ (3.12)
@@
$\frac{T_{ev}}{M_{0}}=\frac{M_{0}^{2}}{3\hbar H_{M}\left(\frac{Q}{M}\right)}=\frac{1}{\varepsilon_{0}}$ (3.13)
3.2 Choice of initial mass parameter
Since we wish to cover significant evaporation in our numerical simulation (we typically manage to evaporate about 53% of the initial mass and charge), our numerical grid must cover a grid length of order $\frac{T_{ev}}{M_{0}}\left(\frac{du}{M_{0}}\right)^{-1}$points. ($du$ is the size of a single grid step, and we usually have $\frac{du}{M_{0}}=\frac{1}{100}$.) For astrophysical black holes of 10 solar masses, $\frac{T_{ev}}{M_{0}}\approx 10^{82}$, which is impossible to implement numerically. We can implement numerically $\frac{T_{ev}}{M_{0}}=10^{3}$. This value definitely does not correspond to a realistic astrophysical black hole, yet this value is still $\gg 1$. We’ll therefore choose our $M_{0}$ such that $\frac{T_{ev}}{M_{0}}=10^{3}$. The corresponding value of $\varepsilon_{0}$ is $\frac{1}{1000}$.
3.3 Choice of units
@@
Recall the parameter $L_{0}$ from Eq. (2.1), which can be chosen freely. We choose $L_{0}=M_{0}$. In our choice of units, we therefore get $L_{0}=1$.
3.4 Basic scaling of semiclassical $T_{\alpha}^{\;\beta}$
@@
$T_{\alpha}^{\;\beta(SC)}=\frac{\hbar}{M^{2}}\hat{T}_{\alpha}^{\;\beta}$ (3.14)
@@
$T_{\alpha}^{\;\beta(SC)}=\frac{\hbar}{M^{4}}H_{T_{\alpha}^{\;\beta}}\left(\frac{r}{M},\frac{Q}{M}\right)$ (3.15)
@@
$T_{vv}\equiv T_{vv}^{(SC)}=\frac{\hbar}{M^{4}}H_{T_{vv}}\left(\frac{r}{M},\frac{Q}{M}\right)$ (3.16)
@@
In Eq. (2.42) we defined the quantity $\widetilde{T}_{vv}\equiv 8\pi r^{2}T_{vv}$. Correspondingly we define $\widetilde{T}_{vv}^{(+)}$ , $\widetilde{T}_{vv}^{(-)}$ , which are $\widetilde{T}_{vv}$ at the two horizons, EH and IH respectively. They scale according to:
$\widetilde{T}_{vv}^{(+)}=8\pi r_{+}^{2}\frac{\hbar}{M^{4}}H_{T_{vv}}\left(\frac{r_{+}}{M},\frac{Q}{M}\right)\equiv\frac{\hbar}{M^{2}}H_{T_{vv}}^{(+)}\left(\frac{Q}{M}\right)$ (3.17)
$\widetilde{T}_{vv}^{(-)}=8\pi r_{-}^{2}\frac{\hbar}{M^{4}}H_{T_{vv}}\left(\frac{r_{-}}{M},\frac{Q}{M}\right)\equiv\frac{\hbar}{M^{2}}H_{T_{vv}}^{(-)}\left(\frac{Q}{M}\right)$ (3.18)
3.5 The charge function $Q(u,v)$
We would like the charge of the black hole to evaporate as well as the mass, so that the mass could evaporate completely. (If the charge does not evaporate, the mass will stop evaporating when the black hole approaches extremality. We are interested in the case of significantly evaporating mass.) If there is a charged quantum field, then $Q$ will indeed evaporate, and we will have electric currents in spacetime. We will therefore assume that there is such a charged quantum field. Hence $Q$ is no longer a constant parameter, but a function $Q(u,v)$ of $u$ and $v$. At the horizon, we have
$Q=Q^{v}(v)$ (3.19)
@@ 3.5 The charge function $Q(u,v)$
$Q(u,v)=Q^{v}(v)+\delta Q(u,v)$ (3.20)
Note that from Eq. (2.27) if $J_{u}=0$ throughout, then $Q_{,u}=0$, and hence $\delta Q(u,v)=0$ because $\delta Q$ vanishes at the EH.
@@
$\frac{Q^{v}(v)}{M(v)}=const$ (3.21)
Note that we choose $0<\frac{Q^{v}}{M}<1$. The derivation of the analytical approximation below in section 4 is not valid in the extremal case because this approximation is based on the assumption $\kappa_{-}>0$.
To what extent is this choice realistic? Let us consider the (arguably hypothetical  [^4] This case is not regarded as realistic because the plasma that typically surrounds black holes would discharge them very quickly.) case of an astrophysically large black hole which is significantly charged initially. Such a black hole would lose charge due to the Schwinger pair production effect, where the strong electric field near the EH causes electron-positron pairs to be created from the vacuum: one of the charged particles would fall into the black hole, reducing its charge, and the other would escape to infinity. This effect, if possible, would quickly discharge the black hole, especially since the charge-to-mass ratio for electrons is $\approx 2\cdot 10^{21}$. A detailed analysis of schwinger pair production can be found in Ref [29]. Schwinger pair production is effective only if the electric field at the event horizon exceeds a critical value $E_{cr}$. The electric field at the EH near extremality is of order $E_{eh}=\frac{Q}{r_{+}^{2}}\approx\frac{Q}{M^{2}}=\frac{Q}{M}\frac{1}{M}$. Therefore, if $\frac{Q}{M}\frac{1}{M}>E_{cr}$, Schwinger pair production will quickly discharge the black hole until $\frac{Q}{M}\frac{1}{M}=E_{cr}$. This means there is a critical mass $M_{cr}=\frac{1}{E_{cr}}$, such that a black hole with $M>M_{cr}$ will discharge very slowly (as $E_{eh}<E_{cr}$). In this range the black hole would lose its mass via Hawking radiation until it reaches extremality. On the other hand, a black hole with $M<M_{cr}$ will discharge while approximately keeping $E_{eh}=E_{cr}$, that is, $\frac{Q}{M}\sim ME_{cr}$. This means that in this range, $\frac{Q}{M}$ decreases roughly linearly with $M$. This varying $\frac{Q}{M}$ would result in a more complicated dynamics inside the evaporating black hole, which we will here avoid for the sake of simplicity.
@@
$\delta Q=\log\left(1+e^{\frac{u+v}{M(v)}}\right)M(v)^{-1}\varepsilon_{0}M_{0}^{2}\alpha$ (3.22)
This structure of $\delta Q$ contributes currents flowing in the $r_{*}=\frac{u+v}{2}$ direction (at leading order in $\varepsilon_{0}$). $\alpha$ is a free dimensionless parameter controlling the strength of the semiclassical currents, which we typically choose to be of order unity. For the dependence of $\delta Q$ on $u+v$, we chose it to be via the function $f(x)=\log\left(1+e^{x}\right)$ as it interpolates smoothly between $f(x)=0$ at $x\ll 0$ (as desired at the EH) and $f(x)=x$ at $x\gg 0$ (corresponding to the region near the IH). The argument of this function $f(x)$ is $x=\frac{u+v}{M(v)}$, which is a dimensionless variable proportional to $r_{*}=\frac{u+v}{2}$. Since $\delta Q$ is initiated from semiclassical currents, it should be proportional to $\hbar$ ($\propto\varepsilon_{0}M_{0}^{2}$). The later has mass squared dimensions, and the dimensions of $Q$ is $M^{1}$, therefore we also need to multiply the expression by $M(v)^{-1}$ to get the correct mass dimensions.
Note that we use $M(v)$ to get the correct mass dimensions, and not $M_{0}$. Because the drift in the mass is extremely slow, at every point we have a very good local RN approximation. We therefore may naturally expect that the semiclassical currents – and therefore also $\delta Q$ – will be well approximated by the corresponding semiclassical quantities on RN background. From dimensional arguments, semiclassical $\delta Q$ on RN background should take the form $\hbar M^{-1}H_{\delta Q}(\frac{r}{M},\frac{Q}{M})$. We assume that this form also holds in the evaporation case, as explained above, and we also substitute $\hbar=\left[3H_{M}\left(\frac{Q}{M}\right)^{-1}\right]\varepsilon_{0}M_{0}^{2}$. For reasons explained above, we chose $H_{\delta Q}$ to be of the form corresponding to Eq. (3.22). It should be emphasized that $M_{0}$ only appears (squared) along with $\varepsilon_{0}$ as a replacement for $\hbar$.
The main motivation for introducing $\delta Q$ is to demonstrate numerically that semiclassical drift in the charge $Q$ that may take place and even become significant near the IH does not affect the backreaction of the metric near the IH – once the four input functions $M(v)$, $Q^{v}(v)$, $\widetilde{T}_{uu}^{(-)}(v)$, $\text{$\widetilde{T}_{vv}^{(-)}$}(v)$ have been specified. (Note that $\delta Q$ may have an effect on the value of $\widetilde{T}_{uu}^{(-)}(v)$, $\text{$\widetilde{T}_{vv}^{(-)}$}(v)$, similar to that of the sources $Z_{R}$$Z_{S}$.) This fact is already embodied in the analytical approximation – where the charge function required is $Q^{v}(v)$ only. Note that in the absence of $\delta Q$, from Eq. (2.27) $J_{u}=0$. Non-vanishing $\delta Q$ allows us to explore scenarios with $J_{u}$ as well. Note that $\delta Q$ is just the simplest model we could choose which will model current $J_{u}$.
3.6 Choice of source functions $Z_{R}$, $Z_{S}$
As already mentioned in subsection 2.8, for the task of validating the analytical approximation, we can freely choose the semiclassical functions which are $T_{uv}^{(SC)}$ and $T_{\theta\theta}^{(SC)}$. This amounts to freely choosing the two source functions $Z_{R}$, $Z_{S}$ defined in Eqs. (2.1172.118).
@@ 3.6 Choice of source functions $Z_{R}$, $Z_{S}$
$Z_{R}=\varepsilon_{0}M_{0}^{2}M^{-3}e^{S}L_{0}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)$ (3.23)
$Z_{S}=\varepsilon_{0}M_{0}^{2}M^{-5}e^{S}L_{0}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)$ (3.24)
@@
The reasoning behind this choice is as follows. In the semiclassical theory these source functions vanish both at the EH and the IH. We therefore choose these functions to be proportional to $e^{S}$, which is close to zero near both the EH and IH [^5] For RN background in Unruh state, it was found [30] that near the IH $Z_{R}$ is proportional to $e^{S}\sim e^{-2\kappa_{-}r_{*}}$, whereas $Z_{S}$ decays as $r_{*}^{-5}$. In our numerical simulation, we decided to simplify the computation scheme by assuming both to be proportional to $e^{S}$ near the IH. We do not expect this to change the resultant backreaction metric because $r_{*}^{-5}$ vanishes sufficiently fast.. We then multiply $e^{S}$ by the length scale $L_{0}$ similar to the way it appears in the evolution equations, see Eqs. (2.37,2.38). The source term is semiclassical, therefore it should be proportional to $\hbar$ ($\propto\varepsilon_{0}M_{0}^{2}$) which has mass squared dimensions, so we want to create a dimensionless quantity from that, leading us to the dimensionless factor $\hbar M^{-2}\propto\varepsilon_{0}M_{0}^{2}M^{-2}$. Since $F_{1}(R)]=\left[\frac{1}{\sqrt{R}}\right]=[l]^{-1}$, the $Z_{R}$ source term is overall multiplied by $M^{-3}$, and since $F_{2}(R)]=\left[\frac{1}{R^{\frac{3}{2}}}\right]=[l]^{-3}$, the $Z_{S}$ source term is overall multiplied by $M^{-5}$. For convenience, we chose the spacetime dependence of these sources to be dependent on $\frac{r}{M}$ only, and we further restricted our choice to linear functions of this variable, determined by the four arbitrary parameters $Z_{R0}$$Z_{Rr}$$Z_{S0}$$Z_{Sr}$. When running our numerical code, we’ll choose multiple specific (arbitrary) combinations of these parameters and the charge parameter $\alpha$.
@@
The presence of these source terms, and / or the changing charge, cause non-conservation of $\widetilde{T}_{vv}$ and $\widetilde{T}_{uu}$, as already mentioned in subsection 2.8. We shall now calculate the amount of this non-conservation, using the evolution equations and the source terms in Eqs. (3.23,3.24). We start by taking the derivative of the constraint Eq. (2.43):
$\left(\widetilde{T}_{uu}\right)_{,v}$ $=R_{,uv}S_{,u}+R_{,u}S_{,uv}-\left(R_{,uv}\right)_{,u}$ (3.25)
$=L_{0}e^{S}\left(F_{1}(R,Q)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)S_{,u}$ (3.26)
$+\left(e^{S}L_{0}F_{2}(R,Q)+\varepsilon_{0}M^{-5}e^{S}L_{0}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,u}$ (3.27)
$-\left[L_{0}e^{S}\left(F_{1}(R,Q)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)\right]_{,u}$ (3.28)
$=L_{0}e^{S}\left(F_{1}(R,Q)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)S_{,u}$ (3.29)
$+\left(e^{S}L_{0}F_{2}(R,Q)+\varepsilon_{0}M^{-5}e^{S}L_{0}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,u}$ (3.30)
$-L_{0}e^{S}\left(F_{1}(R,Q)+\varepsilon_{0}M^{-3}\left(Z_{R0}+\frac{r}{M}Z_{Rr}\right)\right)S_{,u}$ (3.31)
$-e^{S}L_{0}\frac{\partial F_{1}(R,Q)}{\partial R}R_{,u}-e^{S}L_{0}\frac{\partial F_{1}(R,Q)}{\partial Q}Q_{,u}-\varepsilon_{0}M^{-3}e^{S}L_{0}\frac{r_{,u}}{M}Z_{Rr}$ (3.32)
$=e^{S}L_{0}\left[\left(\varepsilon_{0}M^{-5}\left(Z_{S0}+\frac{r}{M}Z_{Sr}\right)\right)R_{,u}-\frac{2Q}{R^{\frac{3}{2}}}Q_{,u}-\varepsilon_{0}M^{-3}\frac{r_{,u}}{M}Z_{Rr}\right]$ (3.33)